What is the ΔH° for the following reaction (in kJ/mol of reaction exactly as written)?

CH4 (g) + 2 Cl2 (g) → CH2Cl2 (g) + 2 HCl (g)

To solve, use Appendix II in your textbook (containing standard heats of formation).
1
Expert's answer
2019-10-18T08:05:49-0400

CH4 (g) + 2Cl2 (g) -> CH2Cl2 (g) + 2HCl (g)

∆fHgas (CH2Cl2) = -95.52 kJ/mol

∆fHgas (HCl) = -92.3 kJ/mol

∆fHgas (CH4) = - 74.87 kJ/mol

∆H°= ∆H(CH2Cl2) + 2∆H(HCl) - ∆H(CH4)= -95.52-2*92.3+74.87=-205.25 kJ/mol